OpenStudy (anonymous):

Please help! What is the tenth term of the geometric sequence that has a common ratio of 1/3 and 36 as its fifth term? A: 4/27 B: 27/4 C: 4/81 D: 1/36

OpenStudy (jdoe0001):

\(\bf \large {a_n=a_1\cdot r^{n-1}\qquad a_5=36\quad \textit{common ratio }\cfrac{1}{3}\qquad thus\\ \quad \\ a_5=a_1\cdot \left(\cfrac{1}{3}\right)^{5-1}\implies 36=a_1\cdot \left(\cfrac{1}{3}\right)^{4}}\) find \(\bf a_1\) or the 1st term, keeping in mind that the 10th term is \(\bf \large a_{10}=a_1\cdot \left(\cfrac{1}{3}\right)^{10-1}\)